OpenStudy (anonymous):

Use the quadratic formula to sovle the equation. Round your answer to the nearest tenth. 5x2 + 2x - 9 = 0

OpenStudy (muzzack):

what is the quadratic formula?

OpenStudy (muzzack):

What is the quadratic formula? The quadratic formula is used to solve a very specific type of equation, called a quadratic equation. These equations are usually written in the following form: Ax^2 + Bx + C = 0 @amistre64 right

OpenStudy (muzzack):

Example: Solve this equation for x: x^2 - 4x + 4 = 0 Solution: What are the three coefficients (a,b,c)? Remember that "a" is the coefficient in front of the x^2 term, b is the coefficient in front of x, and c is that constant at the end. Therefore, for this equation, a=1, b= -4, and c= 4. Plug those values into the quadratic formula and solve for x:

OpenStudy (muzzack):

@Hero @ParthKohli @paki @Preetha

OpenStudy (muzzack):

\[x=\frac{ 4 \pm \sqrt{(-4)^{2} - 4(1)(4)} }{ 2(1) }\]

OpenStudy (muzzack):

now you simplify it

OpenStudy (amistre64):

that is what the quadratic formula is used for yes :)

OpenStudy (muzzack):

oh thx @amistre64

OpenStudy (anonymous):

So the answer is 1.2, and -1.6? @amistre64

OpenStudy (amistre64):

5x2 + 2x - 9 = 0 -5 +- sqrt(25-4(-9)(5)) --------------------- 2(5) -5 +- sqrt(25+180) --------------------- 10 -5 +- sqrt(205) --------------- 10